# The linear targeting problem

### v2

@@ 1. Introduction
If $AX=Y$, then $\operatorname{null}X\subseteq\operatorname{null}Y$, so this condition (or something that implies it) is found in all of our results. The invertible linear targeting problem is just the row equivalence problem, which has a solution if and only if $\operatorname{null}X=\operatorname{null}Y$ [1, Theorem 4.4.1].
If $m<n$ and $AX=Y$, then after a simultaneous permutation of the columns of $X$ and $Y$ we may assume that $X=[X_{1}\,\,X_{2}]$, $Y=[Y_{1}\,\,Y_{2}]$, $X_{1},Y_{1}\in\mathsf{M}_{m\times k}(\mathbb{F})$, $k\leq m$, and $\operatorname{col}X_{2}\subseteq\operatorname{col}X_{1}$. For example, we could choose $k=\operatorname{rank}X$, but any $k$ between $\operatorname{rank}X$ and $m$ will do. Since there is some $B\in\mathsf{M}_{k\times(n-k)}(\mathbb{F})$ such that $X_{2}=X_{1}B$ [1, Theorem 1.6.21.a], we have $AX_{1}=Y_{1}$, from which it follows that $AX_{2}=A(X_{1}B)=Y_{1}B=Y_{2}$. Thus, in considering the linear targeting problem $AX=Y$ with $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$, it suffices to consider only the case $m\geq n\geq 1$.
For any $Z\in\mathsf{M}_{m\times n}(\mathbb{F})$ ($m\times n$ matrices over the field $\mathbb{R}$ or $\mathbb{C}$), $Z^{*}$ is the conjugate transpose of $Z$ and $Z^{\dagger}$ is its Moore–Penrose pseudoinverse; see [1, § 17.4] or [2, 7.3.P7]. The Euclidean norm of $\mathbf{x}\in\mathbb{F}^{n}$ is denoted by $\|\mathbf{x}\|=(\mathbf{x}^{*}\mathbf{x})^{1/2}$. If $A\in\mathsf{M}_{n}(\mathbb{F})$, we write $A\geq 0$ (respectively, $A>0$) if $A\in\mathsf{M}_{n}(\mathbb{F})$ is positive semidefinite (respectively, positive definite), that is, $A$ is Hermitian and $\mathbf{u}^{*}A\mathbf{u}\geq 0$ (respectively, $\mathbf{u}^{*}A\mathbf{u}>0)$ for all nonzero $\mathbf{u}\in\mathbb{F}^{n}$. If $B,S\in\mathsf{M}_{m}(\mathbb{F})$ and $S$ is invertible, then $S^{*}BS$ is congruent to $B$. If $B$ is, respectively, invertible, Hermitian, or positive semidefinite, then $S^{*}BS$ has the same respective properties.
@@ 2. Unconstrained linear targeting
The following theorem re-examines the unconstrained linear targeting problem in the context of a singular value decomposition of the data matrix $X$. Its proof introduces notation and relations that we use throughout our discussion. If $X=0$, the linear targeting problem has a solution if and only if $Y=0$, so this case is not interesting.
@@
in which $R\in\mathsf{M}_{r\times(m-r)}(\mathbb{F})$ and $S\in\mathsf{M}_{m-r}(\mathbb{F})$ are arbitrary; they (and $B_{2})$ are absent if $X$ has full rank. Then
@@
If a property $\mathcal{P}$ is invariant under unitary similarity, the preceding theorem shows that the $\mathcal{P}$-targeting problem is equivalent to a matrix completion problem. The source and target matrices determine $B_{1}$; if we can choose $B_{2}$ so that $B$ (and hence also $VBV^{*})$ has property $\mathcal{P}$, then the $\mathcal{P}$-targeting problem has a solution. For example, if $\mathcal{P}$ is the property that $A$ is invertible, then $B$ must be invertible and hence the data must ensure that $B_{1}$ has full rank. The following corollary reveals a familiar criterion for row equivalence of two matrices; see [1, Theorem 4.4.1 and P.8.30].
@@ 3. Two Lemmas
With the notation in the preceding section, the following lemma memorializes some consequences of the basic assumption $\operatorname{null}X\subseteq\operatorname{null}Y$.
@@
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. If $\operatorname{null}X\subseteq\operatorname{null}Y$, then $X^{*}Y$ is unitarily similar to $\Sigma_{r}Z_{1}\oplus 0_{n-r}$, in which $Z_{1}=V_{1}^{*}YW_{1}$.
@@ 4. Hermitian targeting
If a nonzero $Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ is given, what are all the $X\in\mathsf{M}_{m\times n}(\mathbb{F})$ such that $Y=AX$ for some Hermitian $A?$
@@ 5. Positive semidefinite targeting
We now investigate the positive semidefinite and positive definite targeting problems. An important property of positive semidefinite matrices $A$ is that $A\mathbf{u}=\mathbf{0}$ if and only if $\mathbf{u}^{*}A\mathbf{u}=0$; see [1, Corollary 15.1.18] or [2, Observation 7.1.6].
@@
[…]
@@
$\operatorname{null}H=\operatorname{null}B_{1}=\operatorname{null}H\cap\operatorname{null}L=\operatorname{null}H\cap\operatorname{null}(L^{*}L),$
which implies that
@@ 6. Unitary targeting
($\Leftarrow$) Let $r=\operatorname{rank}X$. If $X^{*}X=Y^{*}Y$, then $\operatorname{null}X=\operatorname{null}Y$ [1, Theorem 15.1.9] and (3) ensures that $B=\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}$ satisfies (2) with $B_{1}=Z\Sigma_{r}^{-1}$. It suffices to show that $B_{1}$ has orthonormal columns, since we may then choose $B_{2}$ to make $B$ unitary; see [1, Corollary 6.3.14]. Observe that
@@
The polar decomposition [1, § 16.3] or [2, § 7.3] provides an alternative approach to the unitary targeting problem. If $X^{*}X=Y^{*}Y$ and $Q=(X^{*}X)^{1/2}$, then $X=U_{1}Q$ and $Y=V_{1}Q$, in which $U_{1},V_{1}\in\mathsf{M}_{m\times n}(\mathbb{F})$ have orthonormal columns. If $U=[U_{1}\,\,U_{2}]$ and $V=[V_{1}\,\,V_{2}]$ are unitary, then $VU^{*}$ is unitary and
@@ 8. Orthogonal Projection Targeting
In the preceding theorem, it is not necessary to assume that $\operatorname{null}X\subseteq\operatorname{null}Y$ because the given hypothesis ensures that
$\operatorname{null}X\subseteq\operatorname{null}(Y^{*}X)=\operatorname{null}(Y^{*}Y)=\operatorname{null}Y.$
We can use a singular value decomposition to identify all solutions $X$ to the orthogonal projection targeting problem $AX=Y$ for a given $Y$.
@@ 10. Normal Targeting
For $n\geq 2$, however, it is not (yet) clear how to proceed. For a normal $A\in\mathsf{M}_{m}(\mathbb{F})$, we have $A^{*}\!A=AA^{*}$, which is a system of $m^{2}$ equations in the entries of $A$. The equations corresponding to diagonal entries of $A^{*}\!A$ and $AA^{*}$ say that, for each $i$, the $i$th row and the $i$th column of $A$ must have the same Euclidean norm. For example, if $m=3$, $n=2$,
$Y=\begin{bmatrix}0&2\\ 1&0\\ 0&0\end{bmatrix},\quad X=\begin{bmatrix}1&0\\ 0&1\\ 0&0\end{bmatrix},$
and $AX=Y$, then the first two columns of $A$ are determined:
$A=\begin{bmatrix}0&2&?\\ 1&0&?\\ 0&0&?\end{bmatrix}.$
No matter how the third column is chosen, the first column and first row of $A$ cannot have the same norm, so the normal targeting problem for $X$ and $Y$ has no solution. Although the normal targeting problem remains an open problem, the following theorem provides a solution in a special case that provides an alternative solution to the reflection and orthogonal projection targeting problems.
@@
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{C})$ and let $\lambda_{1},\lambda_{2}\in\mathbb{C}$ be distinct.
@@
Let $P=E(E^{*}E)^{\dagger}E^{*}$ and $Q=F(F^{*}F)^{\dagger}F^{*}$, which are, respectively, orthogonal projections onto the respective orthogonal subspaces $\operatorname{col}E$ and $\operatorname{col}F$. Let $A=\lambda P+\mu Q$, which is normal with spectrum $\{\lambda,\mu\}$ [1, § 14.9]. Then
$AE$ $=(\lambda P+\mu Q)E=\lambda PE+\mu QE=\lambda PE=\lambda E,$
$AF$ $=(\lambda P+\mu Q)F=\lambda PF+\mu QF=\mu QF=\mu F,$
@@
A reflection is a normal matrix with spectrum $\{1,-1\}$.Theorem 10.1 says that there is a reflection $A$ such that $AX=Y$ if and only if $C=(Y+X)^{*}(Y-X)=0$. Observe that $C=(Y^{*}Y-X^{*}Y)+(X^{*}Y-Y^{*}X)$ is the sum of a Hermitian matrix and a skew-Hermitian matrix, and that $C=0$ if and only if both its Hermitian part and its skew-Hermitian part are zero. This is the condition in Theorem 7.1.
Acknowledgement. We thank Michael A. Dritschel for pointing out that [4, Corollary 1] gives a version of Theorem 5.1 that is valid for bounded linear operators on a Hilbert space.

### v3

@@ 1. Introduction
If $AX=Y$, then $\operatorname{null}X\subseteq\operatorname{null}Y$, so this condition (or something that implies it) is found in all of our results. The invertible linear targeting problem is just the row equivalence problem, which has a solution if and only if $\operatorname{null}X=\operatorname{null}Y$ [3, Theorem 4.4.1].
If $m<n$ and $AX=Y$, then after a simultaneous permutation of the columns of $X$ and $Y$ we may assume that $X=[X_{1}\,\,X_{2}]$, $Y=[Y_{1}\,\,Y_{2}]$, $X_{1},Y_{1}\in\mathsf{M}_{m\times k}(\mathbb{F})$, $k\leq m$, and $\operatorname{col}X_{2}\subseteq\operatorname{col}X_{1}$. For example, we could choose $k=\operatorname{rank}X$, but any $k$ between $\operatorname{rank}X$ and $m$ will do. Since there is some $B\in\mathsf{M}_{k\times(n-k)}(\mathbb{F})$ such that $X_{2}=X_{1}B$ [3, Theorem 1.6.21.a], we have $AX_{1}=Y_{1}$, from which it follows that $AX_{2}=A(X_{1}B)=Y_{1}B=Y_{2}$. Thus, in considering the linear targeting problem $AX=Y$ with $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$, it suffices to consider only the case $m\geq n\geq 1$.
For any $Z\in\mathsf{M}_{m\times n}(\mathbb{F})$ ($m\times n$ matrices over the field $\mathbb{R}$ or $\mathbb{C}$), $Z^{*}$ is the conjugate transpose of $Z$ and $Z^{\dagger}$ is its Moore–Penrose pseudoinverse; see [3, § 17.4] or [4, 7.3.P7]. The Euclidean norm of $\mathbf{x}\in\mathbb{F}^{n}$ is denoted by $\|\mathbf{x}\|=(\mathbf{x}^{*}\mathbf{x})^{1/2}$. If $A\in\mathsf{M}_{n}(\mathbb{F})$, we write $A\geq 0$ (respectively, $A>0$) if $A\in\mathsf{M}_{n}(\mathbb{F})$ is positive semidefinite (respectively, positive definite), that is, $A$ is Hermitian and $\mathbf{u}^{*}A\mathbf{u}\geq 0$ (respectively, $\mathbf{u}^{*}A\mathbf{u}>0)$ for all nonzero $\mathbf{u}\in\mathbb{F}^{n}$. If $B,S\in\mathsf{M}_{m}(\mathbb{F})$ and $S$ is invertible, then $S^{*}BS$ is congruent to $B$. If $B$ is, respectively, invertible, Hermitian, or positive semidefinite, then $S^{*}BS$ has the same respective properties.
@@ 2. Unconstrained linear targeting
The following theorem re-examines the unconstrained linear targeting problem in the context of a singular value decomposition of the data matrix $X$. Later, we find it convenient to employ a singular value decomposition of $Y$, with analogous notation. If $X=0$, the linear targeting problem has a solution if and only if $Y=0$, so this case is not interesting.
@@
in which $R\in\mathsf{M}_{r\times(n-r)}(\mathbb{F})$ and $S\in\mathsf{M}_{n-r}(\mathbb{F})$ are arbitrary. Then
@@
If a property $\mathcal{P}$ is invariant under unitary similarity, the preceding theorem shows that the $\mathcal{P}$-targeting problem is equivalent to a matrix completion problem. The source and target matrices determine $B_{1}$; if we can choose $B_{2}$ so that $B$ (and hence also $VBV^{*})$ has property $\mathcal{P}$, then the $\mathcal{P}$-targeting problem has a solution. For example, if $\mathcal{P}$ is the property that $A$ is invertible, then $B$ must be invertible and hence the data must ensure that $B_{1}$ has full rank. The following corollary reveals a familiar criterion for row equivalence of two matrices; see [3, Theorem 4.4.1 and P.8.30].
@@ 3. Two Lemmas
The following lemma memorializes some consequences of the basic assumption $\operatorname{null}X\subseteq\operatorname{null}Y$.
@@
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. Suppose that $\operatorname{null}X\subseteq\operatorname{null}Y$ and adopt the notation used in the proof of Theorem 2.2. Then $X^{*}Y$ is unitarily similar to $\Sigma_{r}Z_{1}\oplus 0_{n-r}$, in which $Z_{1}=V_{1}^{*}YW_{1}$.
@@ 4. Hermitian targeting
If a nonzero $Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ is given, what are all the $X\in\mathsf{M}_{m\times n}(\mathbb{F})$ such that $Y=AX$ for some Hermitian $A?$ To analyze this question we start with a singular value decomposition of $Y$ instead of $X$.
@@ 5. Positive semidefinite targeting
We now investigate the positive semidefinite and positive definite targeting problems. An important property of positive semidefinite matrices $A$ is that $A\mathbf{u}=\mathbf{0}$ if and only if $\mathbf{u}^{*}A\mathbf{u}=0$; see [3, Corollary 15.1.18] or [4, Observation 7.1.6].
@@
($\Rightarrow$) If $AX=Y$ and $A\geq 0$, then $X^{*}Y=X^{*}AX$, which is positive semidefinite. To verify the assertion about null spaces, use the fact that $X^{*}Y\geq 0$ and observe that
@@ 6. Unitary targeting
($\Leftarrow$) Let $r=\operatorname{rank}X$. If $X^{*}X=Y^{*}Y$, then $\operatorname{null}X=\operatorname{null}Y$ [3, Theorem 15.1.9] and (3) ensures that $B=\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}$ satisfies (2) with $B_{1}=Z\Sigma_{r}^{-1}$. It suffices to show that $B_{1}$ has orthonormal columns, since we may then choose $B_{2}$ to make $B$ unitary; see [3, Corollary 6.3.14]. Observe that
@@
The polar decomposition [3, § 16.3] or [4, § 7.3] provides an alternative approach to the unitary targeting problem. If $X^{*}X=Y^{*}Y$ and $Q=(X^{*}X)^{1/2}$, then $X=U_{1}Q$ and $Y=V_{1}Q$, in which $U_{1},V_{1}\in\mathsf{M}_{m\times n}(\mathbb{F})$ have orthonormal columns. If $U=[U_{1}\,\,U_{2}]$ and $V=[V_{1}\,\,V_{2}]$ are unitary, then $VU^{*}$ is unitary and
@@ 8. Orthogonal Projection Targeting
We can use a singular value decomposition of $Y$ (note the change of strategy here) to identify all solutions $X$ to the orthogonal projection targeting problem $AX=Y$ for a given $Y$.
@@ 10. Normal Targeting
For $n\geq 2$, however, it is not (yet) clear how to proceed. For a normal $A\in\mathsf{M}_{m}(\mathbb{F})$, we have $A^{*}\!A=AA^{*}$, which is a system of $m^{2}$ equations in the entries of $A$. The equations corresponding to diagonal entries of $A^{*}\!A$ and $AA^{*}$ say that, for each $i$, the $i$th row and the $i$th column of $A$ have the same Euclidean norm. If we partition $A=\big{[}\begin{smallmatrix}B&D\\ C&E\end{smallmatrix}\big{]}\in\mathsf{M}_{m}(\mathbb{F})$, in which $B\in\mathsf{M}_{k}(\mathbb{F})$, then inspection of the $(1,1)$ block of $A^{*}A=AA^{*}$ reveals that $B^{*}B-BB^{*}+C^{*}C=DD^{*}$. Consequently, for given $B$ and $C$, in order for there to be matrices $D$ and $E$ such that $A$ is normal, it is necessary that
$H(B,C)=B^{*}B-BB^{*}+C^{*}C$ (13)
be positive semidefinite. If $m=2k$, this necessary condition is sufficient for $k=2$ (see [1]), but not for $k\geq 3$. Consider
$B=\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix}\quad\text{and}\quad C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}.$
In this case, $B$ is normal and $H(B,C)=C$ is positive semidefinite. However, there is no choice of $D$ and $E$ that makes $A$ normal because the first row and first column of $A$ cannot have the same Euclidean norm. For a discussion of the condition (13) see [1] and [2].
Although the normal targeting problem remains open, the following theorem provides a solution in a special case that provides an alternative solution to the reflection and orthogonal projection targeting problems.
@@
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{C})$ and let $\lambda,\mu\in\mathbb{C}$ be distinct.
@@
Let $P=E(E^{*}E)^{\dagger}E^{*}$ and $Q=F(F^{*}F)^{\dagger}F^{*}$, which are, respectively, orthogonal projections onto the respective orthogonal subspaces $\operatorname{col}E$ and $\operatorname{col}F$. If $\operatorname{rank}E+\operatorname{rank}F<n$, let $R$ be the orthogonal projection onto $(\operatorname{col}E\oplus\operatorname{col}F)^{\perp}$; otherwise, let $R=0_{m}$. Let $A=\lambda P+\mu Q+\lambda R$, which is normal with spectrum $\{\lambda,\mu\}$ [3, § 14.9]. Then
$AE$ $=(\lambda P+\mu Q+\lambda R)E=\lambda PE+\mu QE+\lambda RE=\lambda PE=\lambda E,$
$AF$ $=(\lambda P+\mu Q+\lambda R)F=\lambda PF+\mu QF+\lambda RF=\mu QF=\mu F,$
@@
A reflection is a normal matrix with spectrum $\{1,-1\}$.Theorem 10.1 says that there is a reflection $A$ such that $AX=Y$ if and only if $C=(Y+X)^{*}(Y-X)=0$. Observe that $C=(Y^{*}Y-X^{*}X)+(X^{*}Y-Y^{*}X)$ is the sum of a Hermitian matrix and a skew-Hermitian matrix, and that $C=0$ if and only if both its Hermitian part and its skew-Hermitian part are zero. This is the condition in Theorem 7.1.
Acknowledgement. We thank Michael A. Dritschel for pointing out that [6, Corollary 1] gives a version of Theorem 5.1 that is valid for bounded linear operators on a Hilbert space. We thank the anonymous referee for detailed comments on the initial draft.